Ellen is downloading labor market data for the most recent month, but her connection is slow. So far this is all she has been able to collect: Unemployment rate 6% Participation rate 62.5% Not in the labor force 63.5 million Find the working-age population, the labor force, the number of employed workers, and the number of unemployed workers.
Answer #1

Unemployment rate = (no. Of unemployed / labor force) = 6% = 0.06

Labor force participation rate = (labor force/working age population) = 62.5% = 0.625

Or, 0.625 = labor force/( labor force + not in labor force)

Or,( 1/0.625 )= (labor force + not in labor force)/labor force

Or, (1/0.625) = 1+(not in labor force/labor force)

Or, 1.6 = 1+(63.5 million /labor force)

Or, (63.5 million/labor force) = 0.6

Or, labor force = 63.5 million/0.6 = 105.8 million

Therefore, working age population = labor force + not in LF = (105.8+63.5) million = 169.3 million

If unemployment rate = 6% , then (unemployed/105.8 million) = 0.06

Or, no. Of unemployed = 0.06 * 105.8 million = 6.35 million

Therefore, number of employed = labor force - unemployed = (105.8 - 6.35) million = 99.45 million